# DMDSpeech: Distilled Diffusion Model Surpassing The Teacher in Zero-shot Speech Synthesis via Direct Metric Optimization

### v1

@@ Abstract
Diffusion models have demonstrated significant potential in speech synthesis tasks, including text-to-speech (TTS) and voice cloning. However, their iterative denoising processes are inefficient and hinder the application of end-to-end optimization with perceptual metrics. In this paper, we propose a novel method of distilling TTS diffusion models with direct end-to-end evaluation metric optimization, achieving state-of-the-art performance. By incorporating Connectionist Temporal Classification (CTC) loss and Speaker Verification (SV) loss, our approach optimizes perceptual evaluation metrics, leading to notable improvements in word error rate and speaker similarity. Our experiments show that DMDSpeech consistently surpasses prior state-of-the-art models in both naturalness and speaker similarity while being significantly faster. Moreover, our synthetic speech has a higher level of voice similarity to the prompt than the ground truth in both human evaluation and objective speaker similarity metric. This work highlights the potential of direct metric optimization in speech synthesis, allowing models to better align with human auditory preferences. The audio samples are available at https://dmdspeech.github.io/.
@@ 1 Introduction
Diffusion-based speech synthesis models have demonstrated superior robustness compared to LM-based approaches (Le et al., 2024; Lee et al., 2024). By generating all speech tokens simultaneously rather than sequentially, diffusion models avoid the error accumulation inherent in autoregressive models and offer faster generation times for longer sentences since their inference speed is not directly proportional to speech length (Ju et al., 2024). However, a significant drawback of diffusion models is their reliance on iterative sampling processes, which can be computationally intensive and time-consuming (Liu et al., 2024). Additionally, unlike LM-based models that directly estimate the likelihood of the output token distribution, diffusion models focus on estimating the score function rather than generating speech tokens end-to-end (Yang et al., 2023). This characteristic makes it inconvenient to optimize directly on the target distribution space using techniques such as perceptual loss (Johnson et al., 2016) or reinforcement learning with human feedback (RLHF) (Bai et al., 2022), as the models do not produce explicit token probabilities that can be easily adjusted.
To address these limitations, diffusion distillation techniques have been proposed to reduce sampling time (Salimans & Ho, 2022; Song et al., 2023; Sauer et al., 2023; Yin et al., 2024b). By distilling the diffusion model into a one-step generator, it becomes possible to access the final output directly and apply metric optimization methods to improve the desired attributes in the synthesized speech.
In this work, we introduce DMDSpeech, a distilled diffusion-based speech synthesis model designed for efficient and high-quality zero-shot speech generation. By employing distribution matching distillation (DMD) (Yin et al., 2024b; a), we transform the diffusion teacher model into a distilled student model that generates speech in 4 steps. The distilled model provides a direct gradient pathway between the noise input and the speech output, enabling end-to-end optimization of any differentiable metrics. We leverage two metrics relevant to zero-shot speech synthesis: speaker similarity via speaker verification (SV) loss and intelligibility and robustness via connectionist temporal classification (CTC) loss. With the SV loss, we enhance the model’s ability to reproduce the unique characteristics of the prompt speaker, obtaining higher similarity to the prompt judged by both human evaluations and speaker verification models. Additionally, the CTC loss helps improve word error rate (WER), enhancing the intelligibility and accuracy of synthesized speech and achieving lower WER than ground truth and other end-to-end baseline models.
@@
  • We present a distilled diffusion model for zero-shot speech synthesis that achieves state-of-the-art performance with significantly reduced inference time.
  • We introduce a new speech synthesis framework for end-to-end optimization of perceptual metrics. We demonstrate that optimizing SV loss and CTC loss leads to improvements in speaker similarity and WER, respectively.
  • We present an in-depth analysis of our model’s ablations of DMD and direct metric optimization, demonstrating the correlations between human evaluations and perceptual metrics, while highlighting the trade-off between sampling speed and diversity (mode shrinkage).
@@ 2 Related Works
Zero-Shot Text-to-Speech Synthesis  Zero-shot TTS generates speech in an unseen speaker’s voice using a small reference sample without additional training. Initial methods relied on speaker embeddings from pre-trained encoders (Casanova et al., 2022; 2021; Wu et al., 2022; Lee et al., 2022) or end-to-end speaker encoders (Li et al., 2024a; Min et al., 2021; Li et al., 2022; Choi et al., 2022), which required extensive feature engineering and struggled with generalization, limiting scalability. More recently, prompt-based methods using in-context learning with reference prompts have scaled models using autoregressive (Shen et al., 2024; Le et al., 2024; Ju et al., 2024; Lee et al., 2024; Yang et al., 2024; Eskimez et al., 2024; Liu et al., 2024) and diffusion frameworks (Jiang et al., 2023b; Wang et al., 2023a; c; Jiang et al., 2023a; Peng et al., 2024; Kim et al., 2024; Chen et al., 2024b; Meng et al., 2024; Yang et al., 2024; Lovelace et al., 2023; Liu et al., 2024). While these models scale well, they suffer from slow inference due to iterative sampling. Our model, DMDSpeech, addresses this by combining the scalability of prompt-based methods with the efficiency of non-iterative sampling. Through distribution matching distillation, we transform a diffusion-based TTS model into a student model capable of generating speech in a few steps, accelerating inference and enabling direct metric optimization for state-of-the-art speaker similarity and speech quality.
Diffusion Distillation  Diffusion models generate high-quality audio but suffer from slow inference due to iterative sampling (Popov et al., 2021). Diffusion distillation accelerates this process by training a student model to efficiently replicate the teacher’s behavior. Previous methods approximated the teacher’s ODE sampling trajectories. ProDiff (Huang et al., 2022) used progressive distillation (Salimans & Ho, 2022) to reduce sampling steps, while CoMoSpeech (Ye et al., 2023) and FlashSpeech (Ye et al., 2024) employed consistency distillation (Song et al., 2023). Rectified flow methods, such as in VoiceFlow (Guo et al., 2024) and ReFlow-TTS (Guan et al., 2024), aimed to accelerate sampling by straightening sampling paths. However, these methods often compromise quality by forcing the student to follow the teacher’s path, which may not suit its reduced capacity. An alternative is distribution matching, either adversarially (Sauer et al., 2023; 2024) or via score function matching (Yin et al., 2024b), aligning the student with the teacher in distribution to maintain quality. However, these methods may reduce diversity as the student model might prioritize high-probability regions of the distribution. Additionally, many distillation techniques require generating noise-data pairs (Sauer et al., 2024; Yin et al., 2024b; Liu et al., 2023), which is computationally expensive. We utilize DMD2 (Yin et al., 2024a), which bypasses the need for pair generation and enhances quality through adversarial training. Interestingly, we find that the reduction in diversity through DMD distillation can also reduce the chance of sampling from low-probability regions of the distributions, which can lead to unwanted artifacts and hallucinations. Consequently, distillation may enhance the human perceived quality, similar to recent findings in autoregressive diffusion models (Liu et al., 2024).
Direct Metric Optimization  Optimizing generative models using perceptual metrics has gained attention recently. MetricGAN (Fu et al., 2019) optimized speech enhancement models using PESQ and STOI as rewards in an adversarial setting. Reinforcement learning from human feedback (RLHF) has also been used to improve speech naturalness by optimizing predicted MOS scores (Zhang et al., 2024; Chen et al., 2024a). However, these approaches are challenging to apply to many state-of-the-art models due to non-differentiable components like duration upsamplers (Li et al., 2024b; Ye et al., 2024) or iterative sampling (Lee et al., 2024; Peng et al., 2024). In contrast, DMDSpeech allows end-to-end differentiable speech generation without iterative processes. We employ speaker verification (SV) loss and connectionist temporal classification (CTC) loss to optimize speaker similarity and text-speech alignment. By directly optimizing these metrics, we significantly improve speaker similarity and intelligibility, aligning synthesized speech more closely with human auditory preferences, marking the first application of direct metric optimization in speech synthesis models.
@@ 3.1 Preliminary: End-to-End Latent Speech Diffusion
Figure 1: Overview of the DMDSpeech framework. The framework consists of inference and three main components for training: (1) Inference (upper left): A few-step distilled generator $G_{\theta}$ synthesizes speech directly from noise, conditioned on the text and speaker prompt (red arrow). (2) Distribution Matching Distillation (upper right): Gradient computation for DMD loss where the student score model $g_{\bm{\psi}}$ matches the teacher score model $f_{\bm{\phi}}$ to align the distribution of student generator $G_{\theta}$ with the teacher distribution (purple arrow). (3) Multi-Modal Adversarial Training (lower right): The discriminator $D$ uses stacked features from the student score model to distinguish between real and synthesized noisy latents conditioned on both text and prompt (yellow arrows). (4) Direct Metric Optimization (lower left): Direct metric optimization for word error rate (WER) via CTC loss (pink arrow) and speaker embedding cosine similarity (SIM) via SV loss (blue arrow).
@@ 4.1 Model Training
Table 1: Comparison between our models and non-E2E baselines on four subjective metrics: naturalness (MOS-N), sound quality (MOS-Q), voice similarity (SMOS-V), and speaking style similarity (SMOS-S). Scoes are presented as means ($\pm$ standard error). One asterisk (*) indicates a statistically significant difference ($p<0.05$) and double asterisk (**) indicates $p<0.01$ compared to DMDSpeech. The best models and those within one standard error of the best are highlighted.
Model MOS-N MOS-Q SMOS-V SMOS-S Ground Truth 4.47 ($\pm$ 0.03) 4.61 ($\pm$ 0.03) 3.86 ($\pm$ 0.05)∗∗ 3.81 ($\pm$ 0.05)∗∗ Ours (DMDSpeech, N=4) 4.42 ($\pm$ 0.03) 4.59 ($\pm$ 0.03) 4.49 ($\pm$ 0.03) 4.30 ($\pm$ 0.03) Ours (Teacher, N=128) 4.32 ($\pm$ 0.04) 4.55 ($\pm$ 0.03) 4.17 ($\pm$ 0.04)∗∗ 4.00 ($\pm$ 0.04)∗∗ NaturalSpeech 3 (Ju et al., 2024) 4.24 ($\pm$ 0.04)∗∗ 4.55 ($\pm$ 0.03) 4.44 ($\pm$ 0.03) 4.25 ($\pm$ 0.04) StyleTTS-ZS (Li et al., 2024b) 4.40 ($\pm$ 0.03) 4.54 ($\pm$ 0.03) 4.34 ($\pm$ 0.04)∗∗ 4.20 ($\pm$ 0.03)
@@
Model MOS-N MOS-Q SMOS-V SMOS-S Ground Truth 4.37 ($\pm$ 0.03) 4.49 ($\pm$ 0.03) 3.51 ($\pm$ 0.05)∗∗ 3.39 ($\pm$ 0.05)∗∗ Ours (DMDSpeech, N=4) 4.27 ($\pm$ 0.03) 4.45 ($\pm$ 0.03) 4.35 ($\pm$ 0.03) 4.16 ($\pm$ 0.03) Ours (Teacher, N=128) 4.22 ($\pm$ 0.04) 4.40 ($\pm$ 0.03) 4.03 ($\pm$ 0.04)∗∗ 3.87 ($\pm$ 0.04)∗∗ DiTTo-TTS (Lee et al., 2024) 4.28 ($\pm$ 0.04) 4.41 ($\pm$ 0.03) 4.16 ($\pm$ 0.04)∗∗ 4.07 ($\pm$ 0.03) VoiceCraft (Peng et al., 2024) 3.76 ($\pm$ 0.05)∗∗ 3.88 ($\pm$ 0.04)∗∗ 3.41 ($\pm$ 0.05)∗∗ 3.37 ($\pm$ 0.05)∗∗ CLaM-TTS (Kim et al., 2024) 3.77 ($\pm$ 0.05)∗∗ 3.87 ($\pm$ 0.04)∗∗ 3.67 ($\pm$ 0.05)∗∗ 3.43 ($\pm$ 0.05)∗∗ XTTS (Casanova et al., 2024) 3.63 ($\pm$ 0.05)∗∗ 3.89 ($\pm$ 0.04)∗∗ 3.25 ($\pm$ 0.05)∗∗ 3.22 ($\pm$ 0.05)∗∗
@@ 4.3 Comparison to Other Models
All baselines, except for NaturalSpeech 3, were evaluated using the 3,711 samples as per Lee et al. (2024). Since we lacked sufficient samples for a direct evaluation of NaturalSpeech 3, its results are taken from their original paper. Table 5 shows that our model achieved the highest speaker similarity score (SIM) to the prompt, even surpassing the ground truth. The Real-Time Factor (RTF) of the distilled model is 13.7 times lower than the teacher model, which is lower than all baseline methods by a large margin. Although our model had a slightly higher WER (1.94) compared to NaturalSpeech 3 (1.81), it is important to note that our model is entirely end-to-end without explicit duration modeling, unlike NaturalSpeech 3. Both DMDSpeech and NaturalSpeech 3 also exhibited lower WER than the ground truth. One point to consider is the high WER of our teacher model, which is mainly due to cutoff at the end of sentences in the training set caused by faulty segmentation with WhisperX. It affects about 10% of the utterances. After distillation, this issue was resolved due to mode shrinkage (discussed in Section 4.4). Moreover, our model demonstrates significantly faster inference speed compared to all baseline models, as it only requires four sampling steps.

### v2

# DMOSpeech: Direct Metric Optimization via Distilled Diffusion Model in Zero-Shot Speech Synthesis

@@ Abstract
Diffusion models have demonstrated significant potential in speech synthesis tasks, including text-to-speech (TTS) and voice cloning. However, their iterative denoising processes are computationally intensive, and previous distillation attempts have shown consistent quality degradation. Moreover, existing TTS approaches are limited by non-differentiable components or iterative sampling that prevent true end-to-end optimization with perceptual metrics. We introduce DMOSpeech, a distilled diffusion-based TTS model that uniquely achieves both faster inference and superior performance compared to its teacher model. By enabling direct gradient pathways to all model components, we demonstrate the first successful end-to-end optimization of differentiable metrics in TTS, incorporating Connectionist Temporal Classification (CTC) loss and Speaker Verification (SV) loss. Our comprehensive experiments, validated through extensive human evaluation, show significant improvements in naturalness, intelligibility, and speaker similarity while reducing inference time by orders of magnitude. This work establishes a new framework for aligning speech synthesis with human auditory preferences through direct metric optimization. The audio samples are available at https://dmospeech.github.io/.
@@ 1 Introduction
As these models scale to handle increasingly diverse speakers and scenarios, ensuring consistent quality and speaker similarity becomes paramount. While directly optimizing relevant perceptual metrics would be a natural approach, it remains a main challenge across all TTS approaches. Traditional models that rely on monotonic alignment and duration predictors (Shen et al., 2024) cannot propagate gradients through these non-differentiable components, preventing true end-to-end optimization of critical elements such as text encoders and duration predictors. While some approaches like YourTTS (Casanova et al., 2022) have attempted to incorporate speaker similarity loss, these architectural limitations resulted in minimal improvements and prevented optimization of other key metrics like word error rate (WER). Modern end-to-end models face different challenges with direct optimization due to their reliance on iterative sampling. Autoregressive models require sampling steps that scale linearly with the length of the generated speech; diffusion models, while more efficient, still require iterative sampling: even the most advanced approaches need at least 16 steps (chen2024f5). This iterative nature not only makes backpropagation computationally prohibitive but also leads to gradient instability. Furthermore, direct optimization through the diffusion process is impossible because intelligible speech can only be generated at low noise levels, making the gradients from perceptual metrics uninstructive at higher noise levels.  [^1] An illustration is available at https://dmospeech.github.io/demo#noise. This suggests that achieving direct metric optimization requires first addressing the fundamental limitations of iterative sampling. While previous work has explored distillation (Salimans & Ho, 2022; Song et al., 2023; Sauer et al., 2023) as a way to reduce sampling steps, these approaches have focused solely on inference speed, consistently showing performance degradation through the distillation process from teacher to student models (bai2023accelerating; Ye et al., 2023). These challenges highlight the need for a fundamentally new approach that not only reduces sampling steps but also enables true end-to-end optimization while maintaining or improving speech quality.
In this work, we introduce Direct Metric Optimization Speech, a distilled diffusion-based speech synthesis model that achieves both superior performance and faster inference compared to existing approaches. Our key innovation is to enable, for the first time, true end-to-end (E2E) optimization of differentiable metrics in TTS, with two technical advances: (1) reducing sampling steps from 128 to 4 via distribution matching distillation (Yin et al., 2024b; a), and (2) providing a direct gradient pathway from the noise input to speech output without non-differentiable components. This allows us to directly optimize speaker similarity and word error rate through speaker verification (SV) and CTC losses respectively, a capability not achievable in previous TTS approaches due to either non-differentiable components like duration predictors (Casanova et al., 2022) or prohibitively expensive backpropagation through hundreds of sampling steps. Our comprehensive experiments demonstrate that this E2E optimization leads to significant improvements across all metrics, outperforming both the teacher model and other recent baselines in both subjective and objective evaluations. Importantly, we discover that our optimized metrics strongly correlate with human perception, and the distillation process induces beneficial mode shrinkage that improves quality in strongly conditional generation by focusing on high-probability regions without compromising output diversity across different prompts and text inputs.
@@
  • We present the first distilled TTS model that consistently outperforms its teacher, achieving superior results while reducing inference time by over 13×.
  • We introduce a framework enabling true E2E optimization of any differentiable metric in TTS, demonstrating its effectiveness through substantial improvements in speaker similarity and word error rate.
  • We provide comprehensive ablations establishing correlations between objective metrics and human perceptions, revealing new insights into sampling speed and diversity trade-offs in strong conditional generation.
@@ 2 Related Works
Zero-Shot Text-to-Speech Synthesis  Zero-shot TTS has evolved significantly in its approach to quality optimization. Early methods relied on speaker embeddings from pre-trained encoders (Casanova et al., 2022; 2021; Wu et al., 2022; Lee et al., 2022) or end-to-end speaker encoders (Li et al., 2024a; Min et al., 2021; Li et al., 2022; Choi et al., 2022), but struggled with generalization and quality optimization due to their reliance on extensive feature engineering and non-differentiable components. More recent prompt-based methods have demonstrated improved scalability using both autoregressive (Shen et al., 2024; Le et al., 2024; Ju et al., 2024; Lee et al., 2024; Yang et al., 2024; Eskimez et al., 2024; Liu et al., 2024) and diffusion frameworks (Jiang et al., 2023b; Wang et al., 2023a; c; Jiang et al., 2023a; Peng et al., 2024; Kim et al., 2024; Chen et al., 2024b; Meng et al., 2024; Yang et al., 2024; Lovelace et al., 2023; Liu et al., 2024). However, these models face fundamental limitations in optimizing perceptual metrics due to their reliance on iterative sampling. DMOSpeech addresses these limitations by enabling true end-to-end metric optimization while maintaining efficient inference.
Diffusion Distillation  Previous approaches to accelerating diffusion models have explored various distillation techniques, each with distinct trade-offs. Progressive distillation (Huang et al., 2022) and consistency distillation (Ye et al., 2023; 2024) attempt to match intermediate states of the teacher’s sampling trajectory, while rectified flow methods (Guo et al., 2024; Guan et al., 2024) focus on straightening these trajectories. However, these approaches often compromise quality by constraining the student to follow the exact path of the teacher, which may be suboptimal for models with reduced capacity. Distribution matching approaches, whether adversarial (Sauer et al., 2023; 2024) or via score function matching (Yin et al., 2024b), offer an alternative by aligning the student with the teacher in distribution rather than trajectory. While these methods typically require computationally expensive noise-data pair generation (Sauer et al., 2024; Yin et al., 2024b; Liu et al., 2023), DMD2 (Yin et al., 2024a) overcomes this limitation by prioritizing efficiency over diversity. Prior studies have typically viewed this tendency to reduce output diversity as a limitation. However, we demonstrate in this work that in strongly conditional generation tasks like TTS, where strict adherence to input text and speaker prompts is required, this diversity reduction can enhance output quality while maintaining sufficient variation across different inputs. This insight makes DMD2 particularly suitable for zero-shot TTS, offering a natural way to balance quality and diversity.
Direct Metric Optimization  While optimizing perceptual metrics has shown promise in speech enhancement through approaches like MetricGAN (Fu et al., 2019) for PESQ and STOI, and recent attempts have explored RLHF for improving naturalness (Zhang et al., 2024; Chen et al., 2024a), implementing these approaches in modern TTS systems has remained challenging. This difficulty stems from architectural limitations such as non-differentiable duration upsamplers (Li et al., 2024b; Ye et al., 2024) or computationally intensive iterative sampling (Lee et al., 2024; Peng et al., 2024). Previous attempts like YourTTS (Casanova et al., 2022) reported minimal improvements from speaker similarity optimization due to their inability to propagate gradients through all model components. DMOSpeech overcomes these limitations by enabling comprehensive optimization of all model elements through direct gradient pathways, marking the first successful demonstration of true end-to-end metric optimization in speech synthesis while maintaining fast inference and high quality.
@@ 3.1 Preliminary: End-to-End Latent Speech Diffusion
Figure 1: Overview of the DMOSpeech framework. The framework consists of inference and three main components for training: (1) Inference (upper left): A few-step distilled generator $G_{\theta}$ synthesizes speech directly from noise, conditioned on the text and speaker prompt (red arrow). (2) Distribution Matching Distillation (upper right): Gradient computation for DMD loss where the student score model $g_{\bm{\psi}}$ matches the teacher score model $f_{\bm{\phi}}$ to align the distribution of student generator $G_{\theta}$ with the teacher distribution (purple arrow). (3) Multi-Modal Adversarial Training (lower right): The discriminator $D$ uses stacked features from the student score model to distinguish between real and synthesized noisy latents conditioned on both text and prompt (yellow arrows). (4) Direct Metric Optimization (lower left): Direct metric optimization for word error rate (WER) via CTC loss (pink arrow) and speaker embedding cosine similarity (SIM) via SV loss (blue arrow).
@@ 4.1 Model Training
Table 1: Comparison between our models and non-E2E baselines on four subjective metrics: naturalness (MOS-N), sound quality (MOS-Q), voice similarity (SMOS-V), and speaking style similarity (SMOS-S). Scoes are presented as means ($\pm$ standard error). One asterisk (*) indicates a statistically significant difference ($p<0.05$) and double asterisk (**) indicates $p<0.01$ compared to DMOSpeech. The best models and those within one standard error of the best are highlighted.
Model MOS-N MOS-Q SMOS-V SMOS-S Ground Truth 4.47 ($\pm$ 0.03) 4.61 ($\pm$ 0.03) 3.86 ($\pm$ 0.05)∗∗ 3.81 ($\pm$ 0.05)∗∗ Ours (DMOSpeech, N=4) 4.42 ($\pm$ 0.03) 4.59 ($\pm$ 0.03) 4.49 ($\pm$ 0.03) 4.30 ($\pm$ 0.03) Ours (Teacher, N=128) 4.32 ($\pm$ 0.04) 4.55 ($\pm$ 0.03) 4.17 ($\pm$ 0.04)∗∗ 4.00 ($\pm$ 0.04)∗∗ NaturalSpeech 3 (Ju et al., 2024) 4.24 ($\pm$ 0.04)∗∗ 4.55 ($\pm$ 0.03) 4.44 ($\pm$ 0.03) 4.25 ($\pm$ 0.04) StyleTTS-ZS (Li et al., 2024b) 4.40 ($\pm$ 0.03) 4.54 ($\pm$ 0.03) 4.34 ($\pm$ 0.04)∗∗ 4.20 ($\pm$ 0.03)
@@
Model MOS-N MOS-Q SMOS-V SMOS-S Ground Truth 4.37 ($\pm$ 0.03) 4.49 ($\pm$ 0.03) 3.51 ($\pm$ 0.05)∗∗ 3.39 ($\pm$ 0.05)∗∗ Ours (DMOSpeech, N=4) 4.27 ($\pm$ 0.03) 4.45 ($\pm$ 0.03) 4.35 ($\pm$ 0.03) 4.16 ($\pm$ 0.03) Ours (Teacher, N=128) 4.22 ($\pm$ 0.04) 4.40 ($\pm$ 0.03) 4.03 ($\pm$ 0.04)∗∗ 3.87 ($\pm$ 0.04)∗∗ DiTTo-TTS (Lee et al., 2024) 4.28 ($\pm$ 0.04) 4.41 ($\pm$ 0.03) 4.16 ($\pm$ 0.04)∗∗ 4.07 ($\pm$ 0.03) VoiceCraft (Peng et al., 2024) 3.76 ($\pm$ 0.05)∗∗ 3.88 ($\pm$ 0.04)∗∗ 3.41 ($\pm$ 0.05)∗∗ 3.37 ($\pm$ 0.05)∗∗ CLaM-TTS (Kim et al., 2024) 3.77 ($\pm$ 0.05)∗∗ 3.87 ($\pm$ 0.04)∗∗ 3.67 ($\pm$ 0.05)∗∗ 3.43 ($\pm$ 0.05)∗∗ XTTS (Casanova et al., 2024) 3.63 ($\pm$ 0.05)∗∗ 3.89 ($\pm$ 0.04)∗∗ 3.25 ($\pm$ 0.05)∗∗ 3.22 ($\pm$ 0.05)∗∗
@@ 4.3 Comparison to Other Models
All baselines, except for NaturalSpeech 3, were evaluated using the 3,711 samples as per Lee et al. (2024). Since we lacked sufficient samples for a direct evaluation of NaturalSpeech 3, its results are taken from their original paper. Table 5 shows that our model achieved the highest speaker similarity score (SIM) to the prompt, even surpassing the ground truth. The Real-Time Factor (RTF) of the distilled model is 13.7 times lower than the teacher model, which is lower than all baseline methods by a large margin. Although our model had a slightly higher WER (1.94) compared to NaturalSpeech 3 (1.81), it is important to note that our model is entirely end-to-end without explicit duration modeling, unlike NaturalSpeech 3. Both DMOSpeech and NaturalSpeech 3 also exhibited lower WER than the ground truth. One point to consider is the high WER of our teacher model, which is mainly due to cutoff at the end of sentences in the training set caused by faulty segmentation with WhisperX. It affects about 10% of the utterances. After distillation, this issue was resolved due to mode shrinkage (discussed in Section 4.4). Moreover, our model demonstrates significantly faster inference speed compared to all baseline models, as it only requires four sampling steps.